OpenStudy (anonymous):

I have started solving this problem but I got stuck at this part. 4x^3-6x^2-4x-5

OpenStudy (anonymous):

Are we just simplifying?

OpenStudy (anonymous):

No, solving

OpenStudy (anonymous):

And checking for extraneous solutions.

OpenStudy (anonymous):

This is an expression, not an equation. It cannot be solved. If you will recheck your question, we can help you.

OpenStudy (anonymous):

Yeah. If we're solving, there has to be an equal sign.

OpenStudy (anonymous):

The original question was 10 over x squared +2x plus 4 over x= 5 over x-2

OpenStudy (anonymous):

Is this the original question? \[\dfrac{10}{x^2+4}=\dfrac{5}{x-2}\]

OpenStudy (anonymous):

You're missing the +2x after the x^2, @gypsy1274

OpenStudy (anonymous):

No, 10/xsquared -2 plus 4/x

OpenStudy (anonymous):

The last part is right

OpenStudy (anonymous):

How's this? \[\dfrac{10}{x^2-2}+\dfrac{4}{x}=\dfrac{5}{x-2}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Was you first step to find a common denominator? If so, what was it?

OpenStudy (anonymous):

No, I multiplied each numerator by the opposite denominator on the left side

OpenStudy (anonymous):

do you mean that you cross multiplied?

OpenStudy (anonymous):

That is not possible since there are 2 terms on the left side.

OpenStudy (anonymous):

Yes that's what I did

OpenStudy (anonymous):

Here is what I did: Multiplied the entire equation by x. That will eliminate the fraction in the second term.

OpenStudy (anonymous):

Then I multiplied the equation by x-2, to eliminate the fraction in the third term.

OpenStudy (anonymous):

Now, I am looking at canceling....

OpenStudy (anonymous):

Canceling isn't working. So I am multiplying by \(x^2-2\).

OpenStudy (anonymous):

The factoring a third degree polynomial.

OpenStudy (anonymous):

In am lost. I am stuck at 10+4/x-2=5/x

OpenStudy (anonymous):

How did you get there?

OpenStudy (anonymous):

If you were following my steps, lets go through them one at a time. What did you get after multiplying the entire equation by x?

OpenStudy (anonymous):

I got 10/x-2 plus 4=5/-2

OpenStudy (anonymous):

You cannot cancel out part of a term, It is an all or nothing deal. So... \[\dfrac{10}{x^2-2} \times x = \dfrac{10x}{x^2-2}\] Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OK. Fix the rest of the equation, let me know what you get and we can move on from there.

OpenStudy (anonymous):

Now I 10x/xsquared-2x+4=5/-2

OpenStudy (anonymous):

Look again at your last term....I think something may be missing...

OpenStudy (anonymous):

Oh, yeah. X-2

OpenStudy (anonymous):

More than that.... \(\dfrac{5}{x-2} \times x=\)?

OpenStudy (anonymous):

5x

OpenStudy (anonymous):

\(\dfrac{5x}{x-2}\)

OpenStudy (anonymous):

Now, multiply through by one of the remaining denominators. Show me your work for each term please.

OpenStudy (anonymous):

Now I am at 10xsquared-2 over xsquared-2x plus 4x-8=5x

OpenStudy (anonymous):

Which denominator did you multiply by?

OpenStudy (anonymous):

X-2

OpenStudy (anonymous):

Stupid moment, don't mind me.

OpenStudy (anonymous):

3rd

OpenStudy (anonymous):

Remember to distribute....The 10x needs to multiply by each term.

OpenStudy (anonymous):

Don't get it

OpenStudy (anonymous):

You multiplied the equation by x-2. The numerator of the first term is 10x so you need to multiply 10x(x-2). \((10x \times x)+ (10x \times -2)\)

OpenStudy (anonymous):

Okay, so then it's 20x-2x?

OpenStudy (anonymous):

Not quite... \(10x \times x=\)?

OpenStudy (anonymous):

20xsquared-2x?

OpenStudy (anonymous):

No, try again.

OpenStudy (anonymous):

Start with \(10x \times x=\)

OpenStudy (anonymous):

10xsquared

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Yes. and \(-2 \times x = \)?

OpenStudy (anonymous):

10x^2 and -2x

OpenStudy (anonymous):

OK. It's getting to be time for me to sign off. \(10x(x-2) = 10x^2 -20x\)

OpenStudy (anonymous):

Put it all together and you get: \[x^3+2x^2-18x+16=0\] Hopefully you will get the same thing when you combine like terms. Unfortunately, at this point, I have to abandon you because I don't know how to factor a third degree polynomial. I suggest you close this question and ask a new one. Just cut and paste from below. `\[x^3+2x^2-18x+16=0\]` And ask someone how to factor it.

OpenStudy (anonymous):

Okay, thankyou so much